OpenStudy (anonymous):

line contains points (2, -2) and (-4, 1). Another line graphed in the same coordinate plane contains points (1, 2) and (-3, 4). Based on the slopes of these lines, are they parallel, perpendicular or neither? parallel perpendicular neither

OpenStudy (anonymous):

m1= (1+2)/(-4-2)=-0.5 m2=(4-2)/(-3-1)=-0.5 The slopes are the same therefore they are parallel

OpenStudy (anonymous):

thank you!!!